SIEGEL MODULAR FORMS AND THETA SERIES

ATTACHED TO QUATERNION ALGEBRAS II

Siegfried Böcherer Rainer Schulze-Pillot

Abstract. We continue our study of Yoshida’s lifting, which associates to a pair of automorphic forms on the adelic multiplicative group of a quaternion algebra a Siegel modular form of degree 2. We consider here the case that the automorphic forms on the quaternion algebra correspond to modular forms of arbitrary even weights and square free levels; in particular we obtain a construction of Siegel modular forms of weight 3 attached to a pair of elliptic modular forms of weights 2 and 4.

Typeset by AMS-TEX